OpenStudy (anonymous):

Solve: b^2+b-12=0. I came up with -4,3

OpenStudy (anonymous):

you can substitute the values you got back into the equation and see if that works

OpenStudy (anonymous):

(-4)^2 +(3)-12=0 16+3-12=0 19-12=17 Is this correct?

OpenStudy (anonymous):

no you would have to substitute two different times - -4 and 3 both are answers, but the equation has two different solutions

jimthompson5910 (jim_thompson5910):

Checking b = -4 b^2+b-12=0 (-4)^2+(-4)-12=0 16 - 4 - 12 = 0 12 - 12 = 0 0 = 0 ... so the value b = -4 is a solution

OpenStudy (anonymous):

so the right way of verifying your solution would be: (-4)^2 -4 -12 = 0 check this and do the same for 3

jimthompson5910 (jim_thompson5910):

Do the same with the other value to confirm

OpenStudy (anonymous):

3^2+3-12=0 9+3-12=0 12-12=0 Is this right?

jimthompson5910 (jim_thompson5910):

good, so you've confirmed that b = 3 is also a solution

OpenStudy (anonymous):

So the answer is -4,3 right? Or do I say b=-4, b=3

jimthompson5910 (jim_thompson5910):

you would say the solutions are b = -4 or b = 3

jimthompson5910 (jim_thompson5910):

Some books use the idea of the solution set, which means you'd write {-4,3}. But the first method works just fine.

OpenStudy (anonymous):

I think my book goes by solution set...You guys are awesome;-) Thanks!!!

jimthompson5910 (jim_thompson5910):

you're welcome